OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x - 2y = 8.

OpenStudy (anonymous):

and thse ar the answer choices a. x intercept = (8, 0); y intercept = (0, 4) b. x intercept = (8, 0); y intercept = (0, -4) c. x intercept = (-8, 0); y intercept = (0, -4) d. x intercept = (-2, 0); y intercept = (0, 8)

OpenStudy (anonymous):

so can u help?

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

y = .5x - 4 -4 is the y intercept and the x intercept is...

OpenStudy (anonymous):

5

OpenStudy (anonymous):

no I'm figuring it out now :D

OpenStudy (anonymous):

8 is the x intercept

OpenStudy (anonymous):

hey ill brb please stay on my mom just needs me real quick

OpenStudy (anonymous):

B

OpenStudy (anonymous):

ok im bak

OpenStudy (anonymous):

B's the correct answer. I suggest you look up y = mx + b

OpenStudy (anonymous):

ok thanks ima inbox u a question

OpenStudy (anonymous):

and ax+by=c

OpenStudy (anonymous):

i haven learned tht yet